OpenStudy (selena2345):

The temperature in a room is 68 degrees Fahrenheit. What is the temperature in Celsius ?

myininaya (myininaya):

Fahrenheit to Celsius (°F - 32) x 5/9 = °C Celsius to Fahrenheit (°C × 9/5) + 32 = °F

myininaya (myininaya):

\[((68-32) \cdot \frac{5}{9})^o=C\]

myininaya (myininaya):

Where there was F I just plugged in 68

myininaya (myininaya):

you got it selena?

myininaya (myininaya):

let me know if you are having trouble with something

OpenStudy (selena2345):

yes i am having trouble

myininaya (myininaya):

which part? replace F with 68?

OpenStudy (selena2345):

what is the Celsius ?

myininaya (myininaya):

\[C=((68-32) \cdot \frac{5}{9})^o\] that is ^ |

myininaya (myininaya):

All you have to do is simplify using order of operations

myininaya (myininaya):

68-32?

OpenStudy (selena2345):

36

myininaya (myininaya):

now multiply by 5/9

myininaya (myininaya):

\[C=(36 \cdot \frac{5}{9})^o=(4 \cdot 5)^o \text{ since } \frac{36}{9}=4 \]

myininaya (myininaya):

I leave the last operation to you

myininaya (myininaya):

4*5

OpenStudy (selena2345):

The Celsius is 20

myininaya (myininaya):

and put your let degree sign say \[20^o Celsius \]

OpenStudy (selena2345):

thank you myiniaya

myininaya (myininaya):

The temperature in Celsius is 20 degrees if the temperature in Fahrenheit is 68 degrees. :)

OpenStudy (selena2345):

thank you myiniaya